OpenStudy (anonymous):

36 is what % of 450?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

36*100 / 450 = 8

OpenStudy (dumbcow):

"is" means equals "of" means multiplication % implies its a #/100 \[36 = \frac{x}{100}*450\]

OpenStudy (anonymous):

Can you draw it out? I know the process but it wasnt coming out right for me.

OpenStudy (anonymous):

450 - 100% 36 - x x=8 % understand?

OpenStudy (anonymous):

?

OpenStudy (anonymous):



OpenStudy (dumbcow):